In a word processing document or on a separate piece of paper, use the guide to construct a two column proof proving that lines l and m are parallel. Angles 1 and 2 are supplementary by definition. Submit the entire proof to your instructor.
Given:
<1 and <2 are supplementary angles
Prove:
l || m
STATEMENT REASON
1.∠1 and ∠2 are supplementary angles 1. Given
2. m∠1 + m∠2 = 180° 2. ( needs answer)
3. ∠1 and ∠3 are supplementary angles 3. Exterior sides in opposite rays
4. (needs answer) 4. (needs answer)
5. m∠1 + m∠2 = m∠1 + m∠3 5. ( needs answer)
6. ( needs answer) 6. (needs answer)
7. l || m 7. (needs answer)
I just need the blanks filled in

Answer by MathLover1(20850)


STATEMENT ...................................................REASON
1.∠1 and ∠2 are supplementary angles........... 1. Given
2. m∠1 + m∠2 = 180° ........................... 2. Definition of Supplementary Angles
3. ∠1 and ∠3 are supplementary angles.......... 3. Exterior sides in opposite rays
4. m∠1 + m∠3 = 180°.......................................4. (Angles forming a linear pair sum to 180° )
5. m∠1 + m∠2 = m∠1 + m∠3........................ 5. ( Transitive Property of Equality)
6. ( m∠2 =m∠3)..............................................6. (Algebra)
7. l || m ............................................................7. (Converse of Corresponding Angles Postulate)

note:
CONVERSE of the CORRESPONDING ANGLES POSTULATE-
if two lines and a transversal form corresponding angles that are congruent, then the two lines are parallel